In 1995 the life expectancy of a male in a certain country was 62.6 years In 2001 it was 64.9 years Let E represent the life expectancy in year t and let t represent the number of years since 1995
The linear function E(t) that fits the data is
E(t)= t +
(round to the nearest tenth)
Use the function to predict the life expectancy of males in 2005
E(14)=
(round to the nearest tenth)
: In 1995 the life expectancy of a male in a certain country was 62.6 years In 2001 it was 64.9 years Let E represent the life expectancy in year t and let t represent the number of years since 1995
The linear function E(t) that fits the data is
E(t)= t +
(round to the nearest tenth)
Use the function to predict the life expectancy of males in 2005
E(14)=
(round to the nearest tenth)

Answer by stanbon(18749)

In 1995 the life expectancy of a male in a certain country was 62.6 years In 2001 it was 64.9 years Let E represent the life expectancy in year t and let t represent the number of years since 1995
----------
You have two points: (0,62.6) and (6,64.9), where x is the number of years after
1995.
-------------------------------
slope: (64.9-62.6)/(6-0) = 2.3/6 = 23/60
-------------
intercept: when x = 0, y = 62.6
-----------------------------
Equation:
E(t) = (23/60)t + 62.6
---------------------------
The linear function E(t) that fits the data is
(round to the nearest tenth)
Use the function to predict the life expectancy of males in 2005
E(14)= (23/60)*14 + 62.6
E(14) = 68.0
(round to the nearest tenth)
---------------
Cheers,
Stan H.